10cm^3 of a normal specimen of human blood contains 1.2 g of hemoglobin. How
many grams does 31 cm^3 of the same blood cantain?
31 cm^3 of the blood cantain _____ grams of hemoglobin

Answer by stanbon(75887)

10cm^3 of a normal specimen of human blood contains 1.2 g of hemoglobin.
How many grams does 31 cm^3 of the same blood cantain?
-----
Use a proportion:
-----
x/31 = 1.2/10
---
x = 3.1*1.2 = 3.72 g's
-----
Cheers,
Stan H.
============

31 cm^3 of the blood cantain 3.72 grams of hemoglobin